OpenStudy (anonymous):

What is the distance between points P(-3, 5) and Q(3, 1)? Sq root of 13 2 divided by the sq root of 13 3 Divided by the sq root of 13 4 divided by the sq root of 13

OpenStudy (saifoo.khan):

Do you know how to use the distance formula?

OpenStudy (anonymous):

No idea what that is.

OpenStudy (anonymous):

What does that even mean? o.O

OpenStudy (nincompoop):

Definition The distance d between the points \[(x_1,y_1)\] and \[(x_2,y_2)\]is given by, \[d=\sqrt{(x_2-x_1)^{2}+(y_2-y_1)^2}\] Example Problem Given the points ( 1, -2 ) and ( -3, 5 ), find the distance between them. Solution Label the points as follows \[( x_1, y_1 ) = ( -1, -2 )\] and \[( x_2, y_2 ) = ( -3, 5 )\] To find the distance d between the points, use the distance formula : \[d=\sqrt{(-3-1)^2+(5 - (-2))^2}=\sqrt{(-2)^2+(7)^2}=\sqrt{4+49}\] \[d=\sqrt{53}\]

OpenStudy (nincompoop):

@DerickG98` now do you know what it means?

OpenStudy (anonymous):

Nope. Thanks for your help tho, I gave up on that assignment. This does not make any sense to me whatsoever.

OpenStudy (nincompoop):

what part do you not comprehend?

OpenStudy (nincompoop):

do you know how to solve exponents and square roots?

OpenStudy (anonymous):

No

OpenStudy (nincompoop):

then you're not ready for this stuff. maybe you need a remedial course on pre-algebra

OpenStudy (anonymous):

I've already taken remedial courses & pre algebra. Last .5 credit before graduation =(

OpenStudy (nincompoop):

this is what exponents look like \[2\times2=(2)^2\] \[2\times2\times2=(2)^3\] \[3\times3=(3)^2\] \[3\times3\times3=(3)^3\] http://www.mathsisfun.com/exponent.html do you know how to add, subtract, divide and multiply at least?

OpenStudy (anonymous):

Yes i know how to do that :D

OpenStudy (nincompoop):

do you know how to add, subtract, divide and multiply integers?

OpenStudy (anonymous):

Integers are numbers right? Yes.

OpenStudy (nincompoop):

yes they are numbers too http://www.mathsisfun.com/whole-numbers.html

OpenStudy (anonymous):

Yes I know how to do that

OpenStudy (nincompoop):

do you understand the concept of square roots and roots?

OpenStudy (nincompoop):

http://www.mathsisfun.com/algebra/square-root.html

OpenStudy (anonymous):

Sort of.